OpenStudy (anonymous):

Help? Danny has a 15-foot board that he wants to use as a motorcycle ramp. If Danny wants the end of the ramp to be 4 feet off the ground, what angle will be formed between the board and the ground? Round the answer to the nearest whole degree.

OpenStudy (anonymous):

Sin theta = 4/15 Solve for theta

OpenStudy (anonymous):

\[\frac{180 \sin ^{-1}\left(\frac{4}{15}\right)}{\pi }=15.5 {}^{\circ} \]

OpenStudy (anonymous):

15.466 degrees rounds to 15 to the nearest degree.

OpenStudy (anonymous):

Am I wrong? @robtobey

OpenStudy (anonymous):

You are not wrong as far as it goes. In the final analysis @mmschulte wanted the angle between the board and the ground.

OpenStudy (anonymous):

Theta would be the angle between the board and the ground correct?